Two resisters of R1 = 100±3 Ω and R2 = 200±4 Ω are connected on series. Find the maximum absolute error, relative error and percentage error in the equivalent resistance of the combination
1
Expert's answer
2020-11-10T06:59:23-0500

Solution

Maximum absolute error

"\Delta R=\Delta R_1+\Delta R_2"

"\Delta R=3+4=7 \Omega"

Relative error

"\frac{\Delta R}{R}=\frac{\Delta R_1} {R_1}
 +\frac{\Delta R_2} {R_2}"


"\frac{\Delta R}{R}=\frac{3} {100}
 +\frac{4} { 200}=0.07"

Percentage error

"\frac{\Delta R}{R}\times 100=(\frac{\Delta R_1} {R_1}
 +\frac{\Delta R_2} {R_2}) \times100"

Percentage error ="7\%"